OpenStudy (albert0898):

HELP ME understand problems like these. Any tricks, tips, formulas? Thanks! Umberto is a newspaper delivery boy. He delivers newspapers every day of the week. On Sunday, his first day on the job, he delivered newspapers to 18 houses in his neighborhood. Each day afterward, his delivery route included 7 more houses than it did the previous day. How many total newspapers did Umberto deliver in his first 2 weeks on the job?

OpenStudy (anonymous):

on first day, no. of newspaper dilivered= 18 on second day = 18+7= 25 on third day = 25+7= 32 therefore, this forms an AP => 18, 25, 32 ........ where a=18 ; d=7 thus no. of newspaper been delivered in first 2 weeks i.e. 14 days = \[ \frac{ n }{ 2 }*[2a + (n-1)d]\]

OpenStudy (anonymous):

put a= 18, n=14, d=7 and u will get the answer...

OpenStudy (albert0898):

Ah yes! Thanks!